OpenStudy (anonymous):

Harris has a triangular flag. He wants to cut out a similar flag that has dimensions one-fifth of the original flag. If the area of the original flag is 1250 square inches, what is the area of the new flag that Harris cuts out?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

Lets say the base of the original triangle was 5 and the height was 10.1

OpenStudy (anonymous):

not 10.1, 10.0

OpenStudy (anonymous):

the formula of a triangle is 1/2bh

OpenStudy (anonymous):

5x10=50/2=25

OpenStudy (anonymous):

If the dimensions were each multiplied by 1/5, the base would be 1 and the height would be 2.

OpenStudy (anonymous):

1x2/2=1

OpenStudy (anonymous):

Hmm... something's missing...

OpenStudy (anonymous):

Since 1250/25=50 and not 25...

OpenStudy (anonymous):